# Minimax Decision Trees via Martingale Approximations

### v1

@@ 1.2 Regression trees
and the $\mathcal{M}$ in (5) is the class of all piecewise constant functions. The construction procedures of $A_{i}$’ s and $\hat{m}_{N,i}$’ s are data dependent, where the regions are associated to a tree grown sequentially by maximizing the chosen decision criteria, and hence the parameters associated with each tree node are chosen recursively.
The recursive nature of the decision tree also indicates the connection to (P-general). Let $\mathbb{P}_{N}$ denote the empirical measure of the training data and $(\mathbf{X},Y)\buildrel\mathrm{law}\over{\sim}\mathbb{P}_{N}$. A (simplified variant of the) decision tree algorithm produces nested binary partitions $\{\pi_{k}\}_{k\geqslant 0}$ with axis-aligned borders of the input space $\mathbb{R}^{d}$ based on the law of $(\mathbf{X},Y)$, where $\pi_{0}=\{\mathbb{R}^{d}\}$ and the index $k$ is the depth of the tree (Breiman et al., 1984). The prediction $M_{k}$ given by a tree of depth $k$ is defined by (2):
$M_{k}(\omega)=M_{k(N)}(\omega):=\mathbb{E}[Y\mid\mathbf{X}\in A],\quad\text{for }\mathbf{X}(\omega)\in A,\leavevmode\nobreak\ A\in\pi_{k}.$
Note that this notation $M_{k}$ can be interpreted as substituting a random input $\mathbf{X}$ into (7). The estimator of $g_{*}$ is then given by $\hat{m}_{k}(\mathbf{x})=\mathbb{E}[Y\mid\mathbf{X}\in A]$ for $\mathbf{x}\in A,\leavevmode\nobreak\ A\in\pi_{k}$.
To facilitate our discussion, we assume that the number of samples in any $A_{i}\in\pi_{k}$ is nonzero and tends to $\infty$ as $k\rightarrow\infty$ uniformly, and moreover, $N/2^{k}$ is exponentially growing in $k$ [^1] Gordon and Olshen (1984) (Theorem 3.6) imposed a similar set of stronger assumptions that require specific growth rates. Such an assumption will not be required in any of the main results in this paper, but is convenient for the explanation here. In this case, we are interested in obtaining the fast convergence rates of (6), a problem formulated as follows.
(P-regression). Under the model (7) and the above assumption, construct a decision tree algorithm such that if the true signal function $g_{*}$ belongs to a certain wide function class, $\mathbb{E}[\left\lVert\hat{m}_{k}-g_{*}\right\rVert^{2}]$ decays exponentially in $k$.
The classic CART algorithm (Breiman et al. (1984)) for decision trees involves the following widely adopted greedy algorithm that recursively constructs partitions $\{\pi_{k}\}_{k\geqslant 0}$. To partition a set $A\in\pi_{k}$, the greedy algorithm selects a dimension and a threshold where the set $A$ is split into two subsets such that the total remaining risk is minimized. See (11) below for a formal definition. A simplified variant of CART without pruning, called VarianceSplit, is described in Section 3.1.
In the theoretical analysis of the convergence rate of CART in the regression setting, researchers often assume a lower bound on the variance decay when the algorithm splits each node in the decision tree (Chi et al., 2022; Syrgkanis and Zampetakis, 2020; Mazumder and Wang, 2024; Cattaneo et al., 2024). For example, a prevalent variance decay assumption is known as the sufficient impurity decrease (SID) condition introduced by Chi et al. (2022), which requires that each split in the CART algorithm decreases the variance of the parent node by at least a constant factor, leading to exponential decay rates of the MSE of the form $r^{k}$ for some $r\in(0,1)$.
The first general result without the variance decay assumption was obtained by Klusowski and Tian (2024). Their main result asserts that modulo model mis-specification, the MSE of CART is upper bounded by $1/k$ at depth $k$; see (23) below. In a similar vein, Cattaneo et al. (2024) studied the convergence rates of the oblique CART and obtained a $1/k$ upper bound for MSE in the general setting (Theorem 1 therein) as well as an $r^{k}$ upper bound under further assumptions such as node sizes (Theorem 4 therein).
The unconditional rate guarantee of $1/k$ appears significantly slower than the exponential decay rates under variance decay assumptions. On the other hand, Cattaneo et al. (2022) discussed why CART may have slower-than-polynomial convergence. This phenomenon was known as the end-cut preference (ECP) in the literature (see Section 11.8 of Breiman et al. (1984)).
We have chosen to work on the model (7) in the simple regression setting (4), where sequential construction is prevalent (Luo and Pratola, 2023; Luo et al., 2024; Liu et al., 2023). Although there are multiple approaches to construct $\hat{m}_{N}$ in (7) sequentially (Loh, 2014), we focus on the CART-like decision tree model (Breiman et al., 1984) in the regression setting above. The classification setting has been well studied in the computational complexity and encoding literature (Blanc et al., 2020; O’Donnell et al., 2005) in terms of using uniform bounds on Boolean functions between partial trees and optimal trees.
1.3 Our contribution
Partition-based martingale approximations.
[…]

### v2

@@ 2.2 Risk bound
$\mathbb{E}[(M_{k}-M_{k+1})^{2}]=\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2},$ (17)
where for two vertices $v,w$, we write $v\sim w$ if they are connected. Let $\mathcal{F}_{k}$ denote the $\sigma$-algebra generated by the collection ${\mathbbm{1}}_{\{\mathbf{X}\in A\}},\,A\in\pi_{k}$. The nodewise predictor $M_{k}=\mathbb{E}[Y\mid\mathcal{F}_{k}]$ is a Doob martingale in $L^{2}$, and its risk satisfies
$\mathbb{E}\bigl{[}(Y-M_{k})^{2}\bigr{]}=\sum_{i=k}^{\infty}\mathbb{E}\bigl{[}(M_{i+1}-M_{i})^{2}\bigr{]}\quad\text{ and }\quad\mathbb{E}\bigl{[}(M_{k}-M_{k+1})^{2}\bigr{]}=\sum_{A\in\pi_{k}}\sum_{\begin{subarray}{c}A^{\prime}\in\pi_{k+1}\\ A^{\prime}\subset A\end{subarray}}\mathbb{P}(A^{\prime})\bigl{(}\mu_{A^{\prime}}-\mu_{A}\bigr{)}^{2}.$ (18)
In other words, convergence rates reduce to controlling the one-step increments created by each split. The VarianceSplit rule provides no uniform control: under ECP, a child can have arbitrarily small mass, so large within-node fluctuations can be confined to microscopic slivers, and the increment need not contract at a fixed rate. Our MinimaxSplit instead minimizes $\max\{\mathbb{P}(A_{L})\mathrm{Var}(Y\mid A_{L}),\,\mathbb{P}(A_{R})\mathrm{Var}(Y\mid A_{R})\}$, ensuring that at every refinement the worst child risk decreases by a constant factor. Summed along the tree, this yields an exponential contraction of $\mathbb{E}[(M_{k+1}-M_{k})^{2}]$ and hence of $\mathbb{E}[(Y-M_{k})^{2}]$. The cyclic schedule strengthens this guarantee by forcing each coordinate to be refined regularly, preventing all contraction from being spent on a single axis and avoiding symmetry-breaking failures; see Section 2.4.
Figure 3: Visualizing the nested sequence of partitions $\{\pi_{k}\}_{k\geqslant 0}$ and its associated binary subtree representation.
To put the above intuition into play, we apply Hölder’s inequality and (11) to obtain
$\begin{split}\mathbb{E}[(M_{k}-M_{k+1})^{2}]&=\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}\\ &\leqslant\bigg{(}\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}\bigg{)}^{1/3}\\ &\hskip 28.45274pt\times\bigg{(}\max_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}|\ell_{v_{k}}-\ell_{v_{k+1}}|\bigg{)}^{2/3}\\ &\leqslant\mathrm{Var}(Y)^{2/3}2^{-2k/3}\bigg{(}\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}|\ell_{v_{k}}-\ell_{v_{k+1}}|\bigg{)}^{2/3}.\end{split}$ (19)
$\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}|\ell_{v_{k}}-\ell_{v_{k+1}}|\leqslant\sum_{A\in\pi_{k+1}}\Big{(}\sup_{A}g-\inf_{A}g\Big{)}\leqslant\left\lVert g\right\rVert_{\mathrm{TV}}.$ (20)
On the other hand, $\mathrm{Var}(Y)=\mathrm{Var}(g(X))\leqslant\left\lVert g\right\rVert_{\mathrm{TV}}^{2}$. Inserting (20) into (19) and using (18), we obtain $\mathbb{E}[(Y-M_{k})^{2}]\leqslant L2^{-2k/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2}$ for some $L>0$.
Our approach is fundamentally different from Klusowski and Tian (2024) in the following sense. On the one hand, the approach of Klusowski and Tian (2024) relies on aggregating one-step lower bounds on the impurity gain (or the decay of the risk). On the other hand, recall from (12) that the total risk decreases by the amount
$\left(\mathbb{E}[Y\mid A_{L}]-\mathbb{E}[Y\mid A_{R}]\right)^{2}(1-\mathbb{P}(A_{L}))\mathbb{P}(A_{L})$ (21)
after $A$ splits into $A_{L}\cup A_{R}$. The intuition is that although the gain (21) can be small at a certain step, it reveals information on the underlying joint distribution that guarantees that either it has already gone through significant risk decay in previous steps, or it is soon happening, or the total variation $\left\lVert g\right\rVert_{\mathrm{TV}}$ must be large. Let us illustrate the above intuition by a quick example.
Example 5.
Consider the case where $d=1$, $X\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1)$, and $Y=\sin(2\pi 2^{p}X)$. Figure 4 below plots the remaining risk $\mathbb{E}[(Y-M_{k})^{2}],\leavevmode\nobreak\ 1\leqslant k\leqslant K$ with power parameters $p=1,2,3,4$ and maximum tree depth $K=12$. Due to the symmetry of the sine function, the first few splits of the MinimaxSplit do not reduce the remaining risk, since the splitting points are points of symmetry. However, as soon as the splitting breaks through the symmetries, the remaining risk decays exponentially to zero. The VarianceSplit algorithm makes mild progress at every step, but is quickly surpassed by MinimaxSplit once the latter breaks through the symmetries. The reason is that due to ECP, the splitting points are near the boundary of the nodes, leading to unbalanced nodes. In other words, while the MinimaxSplit algorithm may not reduce the risk as steadily as VarianceSplit, the steps with little risk reduction often implicitly overcome certain hard constraints inherent to the regression problem (such as the symmetries of the sine function), paving the way for a much faster risk decay in later stages. Note that the occurrence of minimal risk deduction from the first few steps of MinimaxSplit does not contradict the risk bound from Theorem 2, because the upper bound already incorporates the total variation of the signal function.
Figure 4: Comparison of the mean squared errors for the MinimaxSplit and VarianceSplit regimes, in the setting where $X\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1)$ and $Y=\sin(2\pi 2^{p}X)$ with $p=1,2,3,4$ and maximum tree depth $12$.
2.3 Empirical risk bound for cyclic MinimaxSplit
So far, we have discussed the cases where the joint distribution $(\mathbf{X},Y)$ is marginally atomless, which does not include the class of empirical measures. We now focus on the regression setting and derive finite-sample performance guarantees. Suppose that under the original law $\mathbb{P}_{*}$, $\mathbf{X}_{*}$ is marginally atomless and $Y_{*}\mid\mathbf{X}_{*}=g_{*}(\mathbf{X}_{*})+\varepsilon:=\mathbb{E}[Y_{*}\mid\mathbf{X}_{*}]+\varepsilon,$ where the error $\varepsilon=Y_{*}-g_{*}(\mathbf{X}_{*})$ is sub-Gaussian, i.e., for some $\sigma>0$, $\mathbb{P}_{*}(|\varepsilon|\geqslant u)\leqslant 2\exp(-{u^{2}}/({2\sigma^{2}})),\leavevmode\nobreak\ u\geqslant 0.$ Also, we consider the empirical law of $(\mathbf{X},Y)$ of $N$ samples from $\mathbb{P}_{*}$.
A limitation of Theorem 2 is that it applies only if the covariate $\mathbf{X}$ is marginally atomless (for (10) to hold). Therefore, we need the following result that deals with measures whose marginals may contain atoms. For a random variable $Y$, denote by $\operatorname{supp}Y$ the support of $Y$.
Theorem 6.
Suppose that $\mathbf{X}$ is purely atomic and for some $N>0$,
Let $\Delta Y:=\sup\operatorname{supp}Y-\inf\operatorname{supp}Y$, and let the associated $\{M_{k}\}_{k\geqslant 0}$ be constructed from the cyclic MinimaxSplit algorithm. Then uniformly for any $\delta>0$ and $k\geqslant 0$,
$\begin{split}\mathbb{E}[(Y-M_{k})^{2}]&\leqslant(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\frac{2^{k+2}(\Delta Y)^{2}}{3N}+\inf_{g\in\mathcal{G}}\bigg{(}\Big{(}(1+\delta)+\frac{2(1+\delta)(1+\delta^{-1})}{3\cdot 2^{2\lfloor k/d\rfloor/3}}\Big{)}\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}]\\ &\hskip 170.71652pt+(1+\delta^{-1})\bigg{(}\frac{2}{3}+\Big{(}\frac{1+\delta^{-1}}{4}\Big{)}^{2/3}\bigg{)}2^{-2\lfloor k/d\rfloor/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2}\bigg{)}.\end{split}$ (23)
Our next key result is the following oracle inequality for decision trees under model mis-specification (i.e., when $g_{*}$ does not belong to $\mathcal{G}$). Let $g_{k}$ be the output of the cyclic MinimaxSplit construction (that is, $M_{k}=g_{k}(\mathbf{X})$). Note that $g_{k}$ is random under the law $\mathbb{P}_{*}$. We use $\left\lVert\cdot\right\rVert$ to denote the $L^{2}$ norm in $\mathbb{P}_{*}$, e.g. $\left\lVert g_{k}-g_{*}\right\rVert^{2}=\mathbb{E}[(g_{k}(\mathbf{X}_{*})-g_{*}(\mathbf{X}_{*}))^{2}]$, where the expectation is taken in $\mathbb{P}_{*}$.
Theorem 7.
In addition to the performance and convergence rate guarantees, we briefly comment on the computational efficiency of MinimaxSplit compared to VarianceSplit. Typically, the time complexity consists of sorting and identifying the split dimension and location (Louppe, 2014, Section 5). The time complexity for sorting does not depend on the splitting criteria. In terms of solving for the split dimension and location, our MinimaxSplit algorithm (4) (or its cyclic variant (5)) is advantageous over VarianceSplit (1), because of the monotonicity of the risk as a function of the split location, thus reducing the time complexity from linear in the node size to logarithmic in the node size.
2.4 Why cyclic MinimaxSplit outperforms MinimaxSplit
The key reason why the above risk bounds and the oracle inequality are only derived for the cyclic MinimaxSplit algorithm is that there is an anti-symmetry-breaking preference (ASBP) for MinimaxSplit, which we briefly explain in this section. The gist is that if the underlying model possesses different symmetry conditions across different dimensions, the MinimaxSplit algorithm is likely to always split over the asymmetric dimensions, thus leaving the symmetric dimensions untouched. This leads to an undesirable convergence rate, or even inconsistency in the large-sample regime. We start with a simple example in the marginally atomless case.
Example 8.
Consider $\mathbf{X}=(X_{1},X_{2})$ uniformly distributed on $1,1]^{2}$ and $Y=X_{1}+|X_{2}|$. Clearly, $Y=f(\mathbf{X})$ for some $f\in\mathcal{G}$ with $\left\lVert f\right\rVert_{\mathrm{TV}}=4$. Recall (4) and consider a MinimaxSplit step applied to a splittable set $A=[\alpha,\beta)\times[-1,1]$ where $1\leqslant\alpha<\beta\leqslant 1$. Along the second dimension $x_{2}$, the split location of MinimaxSplit in (4) is always given by $x_{2}=0$, due to the symmetry of the law of $(\mathbf{X},Y)$ along $x_{2}=0$ on the set $A$. However, there is no risk decay given by this split, again due to the above symmetry. On the other hand, along the first dimension $x_{1}$, the split location of MinimaxSplit in (4) is given by $x_{1}=(\alpha+\beta)/2$ and the risk decay is strictly positive. It follows that the $\operatorname*{arg\,min}$ in (4) always satisfies $j=1$. By induction, if we apply the MinimaxSplit algorithm to the joint distribution $(\mathbf{X},Y)$, the split will always be along the first dimension $x_{1}$.
Therefore, if the associated $\{M_{k}\}_{k\geqslant 0}$ is constructed from the MinimaxSplit algorithm, each $M_{k}$ can be written as a function of $X_{1}$ only, say $M_{k}=h_{k}(X_{1})$. But then by the independence of $X_{1}$ and $X_{2}$, for any $k$,
$\begin{split}\mathbb{E}[(Y-M_{k})^{2}]&=\mathbb{E}[(X_{1}-h_{k}(X_{1})+|X_{2}|)^{2}]\\ &=\mathbb{E}[(X_{1}-h_{k}(X_{1}))^{2}]+\mathbb{E}[|X_{2}|^{2}]+2\mathbb{E}[(X_{1}-h_{k}(X_{1}))]\mathbb{E}[|X_{2}|]\\ &\geqslant\mathrm{Var}(X_{1}-h_{k}(X_{1}))+\mathrm{Var}(|X_{2}|)+(\mathbb{E}[X_{1}-h_{k}(X_{1})]+\mathbb{E}[|X_{2}|])^{2}\\ &\geqslant\mathrm{Var}(|X_{2}|)=\frac{1}{12},\end{split}$ (25)
and thus the MSE does not even decay to zero, meaning that the counterpart of (13) cannot hold.
Example 8 does not involve samples, but it is intuitively true that ASBP should still hold if one considers empirical measures of $(\mathbf{X},Y)$ with the goal of estimating $f$. This can be made rigorous using a Donsker class argument, which we discuss in Appendix C. Example 8 also generalizes to higher dimensions, such as using the function $Y=f(\mathbf{X})$ where $f(x_{1},\dots,x_{d})=x_{1}+\dots+x_{d-1}+|x_{d}|$.
This ASBP phenomenon has an important consequence of the inadmissibility of classic random forests in the context of MinimaxSplit. We follow the setting of Section 7 of Klusowski and Tian (2024), which is a slight adaptation of the Breiman’s classic random forest (Breiman, 2001). A random forest consists of a collection of independent subsampled trees, where at each node, the split dimensions are optimized among a random pre-selected set $\mathcal{S}\subseteq[d]$. The set $\mathcal{S}$ is independently randomly generated with a fixed size $m_{\mathrm{try}}\in[d]$ for each node. The notation $m_{\mathrm{try}}$ will be kept throughout this paper. Consider a random forest under the MinimaxSplit regime, with $m_{\mathrm{try}}\geqslant 2$ and $g(\mathbf{x})=x_{1}+\dots+x_{d-1}+|x_{d}|$ on $1,1]^{d}$. Then each $\mathcal{S}$ must include a dimension in [d1][d-1], along which the split always reduces more risk than the $d$-th dimension. This implies that the $d$-th dimension is never reached and the risk is strictly bounded away from zero for any $k$, using an argument similar to (25). Therefore, the consistency of the random forest under the MinimaxSplit regime is not always guaranteed in the large-sample regime (even if the signal is noiseless), unless we pick $m_{\mathrm{try}}=1$. In the next section, we analyze the case $m_{\mathrm{try}}=1$, which we call a random-dimension random forest, and establish the corresponding consistency for the MinimaxSplit random forest. The empirical performance will be presented in Section 5.3.
3 Ensemble models
3.1 MinimaxSplit random forests
The ensemble method we consider here is the original Breiman’s random forest (Breiman et al., 1984), which aggregates multiple decision trees to improve the precision and robustness of the prediction. Each tree is grown to a fixed depth $k$ according to some splitting rule, and the split dimension of each node is optimized among independent selections of random subsets of $[d]$ with a fixed cardinality $m_{\mathrm{try}}$. Each tree in the ensemble is trained on a bootstrap sample of the original dataset and predictions are made by averaging the output of individual trees, effectively reducing the overall variance of the model (see Algorithm 1 in Appendix E). We call the ensemble the MinimaxSplit (resp. VarianceSplit) random forest, if the base learners follow the MinimaxSplit (resp. VarianceSplit) algorithm. [^6] Note that the cyclic MinimaxSplit case does not apply (unless $m_{\mathrm{try}}=1$) due to the lack of an optimization procedure over the dimensions.
3.2 Empirical risk bound for MinimaxSplit random-dimension random forests
Our discussions in Section 2.4 indicate that the random forest approach for the MinimaxSplit algorithms is better performed with $m_{\mathrm{try}}=1$ to guarantee consistency. In other words, we randomize the dimension at every node. As a consequence, applying the MinimaxSplit and the cyclic MinimaxSplit splitting rules yield the same ensemble in distribution. We call the resulting ensemble the MinimaxSplit random-dimension random forest.
Recall the regression setting from Section 2.3 and consider a random forest consisting of trees in which split dimensions at each node are i.i.d. sampled from $[d]$, independently of the data, similar to ExtraTrees (Geurts et al., 2006) but the randomness lies in choosing the dimension instead of the split location. Fix a depth $k$ and the number of samples $N$. Denote by $n$ the number of trees in the forest. Let $\Sigma$ denote the law of the splitting dimensions and $\mathbb{E}_{n}$ denote the expectation with respect to the empirical measure on $n$ samples from $\Sigma$. Let $g_{k}$ be the estimator constructed at deapth $k$ from the MinimaxSplit random-dimension random forest. Our goal is then to bound the $L^{2}$ error $\mathbb{E}\big{[}\left\lVert\mathbb{E}_{n}[g_{k}]-g^{*}\right\rVert^{2}\big{]},$where the outer expectation is taken in terms of randomness both from the empirical samples from $(\mathbf{X}_{*},Y_{*})$ and from the empirical measure $\mathbb{E}_{n}$. We will provide empirical evidence of the effectiveness of such a randomised construction in Section 5.3.
Theorem 9.
Assume the same regression setting of Theorem 7, and the above random forest definition. We have
$\begin{split}\mathbb{E}\big{[}\left\lVert\mathbb{E}_{n}[g_{k}]-g^{*}\right\rVert^{2}\big{]}\leqslant&\frac{C2^{k}(\log N)^{2}\log(Nd)+2^{k}(\log N)dn^{-1/3}}{N}\\ &+2\inf_{g\in\mathcal{G}}\bigg{(}2\left\lVert g-g^{*}\right\rVert^{2}+\frac{8}{3}\Big{(}\left\lVert g-g^{*}\right\rVert^{2}+\left\lVert g\right\rVert_{\mathrm{TV}}^{2}\Big{)}\Big{(}e^{-\frac{k}{6d}}+\frac{d}{n^{1/3}}\Big{)}\bigg{)},\end{split}$ (26)
where $C>0$ is a constant depending only on $\left\lVert g^{*}\right\rVert_{\infty}$ and $\sigma^{2}$. In particular, if $N/(2^{k}(\log N)^{2}\log(Nd))\to\infty$, $k/d\to\infty$, and $n/d^{3}\to\infty$, then the MinimaxSplit random-dimension random forest is consistent.
Let us compare Theorem 9 with the empirical risk bound for a single tree given by Theorem 7. First, the exponent k/(6d)-k/(6d) now does not contain a floor function but has an asymptotically slower rate if $k/d\to\infty$. This is due to the non-cyclicity of the (random) selection of dimensions. Second, the formula (26) does not involve $\delta$ for simplicity, although the same argument works for more general choices of $\delta>0$.
4 Minimax decision trees for classification
Following the C4.5 generalization in Klusowski and Tian (2024), we also provide corresponding results when using MinimaxSplit for entropy loss. Throughout this section, we consider binary response $Y\in\{\pm 1\}$ with covariates $\mathbf{X}\in\mathbb{R}^{d}$ (so we consider $(\mathbf{X},Y)$ as a joint distribution), and $\eta(\mathbf{x})\coloneqq\mathbb{P}(Y=1\mid\mathbf{X}=\mathbf{x})$ denotes the regression function. For any measurable cell $A\subset\mathbb{R}^{d}$ we write
$\eta_{A}\;\coloneqq\;\mathbb{P}(Y=1\mid\mathbf{X}\in A),\qquad w(A)\;\coloneqq\;\mathbb{P}(\mathbf{X}\in A),$
and we measure impurity on $A$ with the Shannon entropy
$H(Y\mid\mathbf{X}\in A)\;=\;h(\eta_{A})\;\coloneqq\;-\eta_{A}\log\eta_{A}-(1-\eta_{A})\log(1-\eta_{A}).$ (27)
Axis-aligned splits refine a cell $A$ into left and right children $A_{L}=\{\mathbf{x}\in A:x_{j}<z\}$ and $A_{R}=A\setminus A_{L}$ for some $j\in[d]$ and $z\in\mathbb{R}$. Standard C4.5 chooses $(j,z)$ to minimize the weighted sum $w(A_{L})h(\eta_{A_{L}})+w(A_{R})h(\eta_{A_{R}})$. In the MinimaxSplit version, we instead choose $(j,z)$ to minimize $\max\{w(A_{L})h(\eta_{A_{L}}),\,w(A_{R})h(\eta_{A_{R}})\}$, which explicitly prevents one child from remaining both large and highly impure. As in the regression setting, when splitting at level $k$, the cyclic MinimaxSplit variant fixes the coordinate at level $k$ to $j_{k}=1+(k\ (\mathrm{mod}\ d))$ and optimizes only over the threshold $z$.
Given a split rule, we obtain a nested sequence of axis-aligned partitions $\{\pi_{k}\}_{k\geqslant 0}$ by applying a single binary split per level. For any $\omega$, we let $A_{k}(\omega)\in\pi_{k}$ be the unique cell containing $\mathbf{X}(\omega)$. The node statistic we track is the log-odds of the average class-probability in that cell,
$M_{k}(\omega)\;\coloneqq\;\log\Big{(}\frac{\eta_{A_{k}(\omega)}}{1-\eta_{A_{k}(\omega)}}\Big{)}.$ (28)
Let $g_{k}(\mathbf{x}):=\log(\eta_{A_{k}}/(1-\eta_{A_{k}}))$ for $\mathbf{x}\in A_{k},\leavevmode\nobreak\ A_{k}\in\pi_{k}$. The predictor associated with depth $k$ is the plug-in classifier $2{\mathbbm{1}}_{\{g_{k}(\mathbf{x})\geqslant 0\}}-1$ (or equivalently, $2{\mathbbm{1}}_{\{\eta_{A_{k}}\geqslant{1}/{2}\}}-1$ for the unique $A_{k}\in\pi_{k}$ that contains $\mathbf{x}$) and the canonical convex surrogate is the conditional cross-entropy on the partition.
With this preparation, the results in this section follow the same logical arc as in the regression setting. Theorem 10 treats the marginally atomless case and shows that, under the cyclic MinimaxSplit schedule, the log-odds process $\{M_{k}\}$ contracts in mean square at an exponential rate that depends only on the dimension and the coordinate range. Theorem 11 covers the complementary purely atomic setting. Here, we need a mild anti-concentration assumption on coordinate masses; the theorem then delivers an explicit upper bound for $\mathbb{E}[(Y-M_{k})^{2}]$ that reflects the discrete resolution $N$ of the atoms. Finally, Theorem 12 translates these bounds into an oracle inequality for the excess classification risk of $g_{k}$ under the 0-1 loss via standard calibration between cross-entropy and misclassification.
Theorem 10.
Suppose that $\mathbf{X}$ is marginally atomless. Let the associated process $\{M_{k}\}_{k\geqslant 0}$ be constructed from (28) using the cyclic MinimaxSplit algorithm. Then uniformly for any $k\geqslant 0$,
$\begin{split}\mathbb{E}[(Y-M_{k})^{2}]&\leqslant\inf_{g\in\mathcal{G}}\Big{(}\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+\sqrt{\frac{\log 2}{4}}\,2^{-\lfloor k/d\rfloor/2}\left\lVert g\right\rVert_{\mathrm{TV}}^{1/2}\Big{)}.\end{split}$ (29)
Theorem 11.
Suppose that $\mathbf{X}$ is purely atomic and for some $N>0$,
Let the associated process $\{M_{k}\}_{k\geqslant 0}$ be constructed from (28) using the cyclic MinimaxSplit algorithm. Then uniformly for any $k\geqslant 0$,
$\begin{split}\mathbb{E}[(Y-M_{k})^{2}]&\leqslant 2^{-\lfloor k/d\rfloor/2}\frac{2^{k}\log N}{N}+\inf_{g\in\mathcal{G}}\Big{(}\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+\sqrt{\frac{\log 2}{4}}\,2^{-\lfloor k/d\rfloor/2}(\left\lVert g\right\rVert_{\mathrm{TV}}^{1/2}+\frac{\left\lVert g\right\rVert_{\mathrm{TV}}}{2})\Big{)}.\end{split}$ (31)
The proofs of Theorems 10 and 11 follow essentially the same principles as in the regression setting, except that now $\{M_{k}\}_{k\geqslant 0}$ no longer forms a martingale. The key is to apply the monotonicity of the map $z\mapsto\mathbb{P}(\mathbf{X}\in A,\,X_{j}<z)\,h(\mathbb{P}(Y=1\mid\mathbf{X}\in A,\,X_{j}<z))$, a fact that mirrors Lemma 2.1 and will be justified in Lemma B.3 below.
In the next result, we state the corresponding oracle inequality. Following Klusowski and Tian (2024), for a map $g$ we consider the misclassification risk $\mathrm{Err}(g):=\mathbb{P}(Y\neq 2{\mathbbm{1}}_{\{g(\mathbf{X})\geqslant 0\}}-1).$ Further, we set $g_{*}(\mathbf{x}):=\log(\mathbb{P}(Y=1\mid\mathbf{X}=\mathbf{x})/\mathbb{P}(Y=-1\mid\mathbf{X}=\mathbf{x}))$.
Theorem 12.
Assume that $\mathbf{X}_{*}$ is marginally atomless. We have that under the cyclic MinimaxSplit construction,
$\begin{split}\mathbb{E}[\mathrm{Err}(g_{k})]-\mathrm{Err}(g_{*})\leqslant L\Big{(}&\frac{2^{k}(\log N)^{2}\log(Nd)}{N}\Big{)}^{1/4}\\ &+\inf_{g\in\mathcal{G}}\Big{(}\left\lVert g-g_{*}\right\rVert^{2}+\Big{(}{\frac{\log 2}{4}}\Big{)}^{1/4}\,2^{1-\lfloor k/d\rfloor/4}(\left\lVert g\right\rVert_{\mathrm{TV}}^{1/4}+\frac{\left\lVert g\right\rVert_{\mathrm{TV}}^{1/2}}{\sqrt{2}})\Big{)},\end{split}$ (32)
where $L>0$ is a positive universal constant.
Remark 13.
The bound (32) can be relaxed to
$\mathbb{E}[\mathrm{Err}(g_{k})]-\mathrm{Err}(g_{*})\leqslant L\Big{(}\frac{2^{k}(\log N)^{2}\log(Nd)}{N}\Big{)}^{1/4}+\inf_{g\in\mathcal{G}}\Big{(}\left\lVert g-g_{*}\right\rVert^{2}+L2^{-\lfloor k/d\rfloor/4}(\left\lVert g\right\rVert_{\mathrm{TV}}+1)\Big{)}$
for some large constant $L>0$, using the rough inequality $x\leqslant\min\{x^{2}+1,x^{4}+1\}$. Compared to the oracle inequality of Klusowski and Tian (2024), we see that we have achieved exponential convergence rates and removed the factor $\left\lVert g\right\rVert_{\infty}$. The proof of Theorem 12 follows in exactly the same way as Theorem 4.3 of Klusowski and Tian (2024) in the C4.5 case, as the arguments therein do not depend on how the tree is constructed, as long as the splitting rule is axis-aligned.
5 Numerical experiments and applications
In this section, we provide experiments that verify the advantages of MinimaxSplit and cyclic MinimaxSplit methods in low-dimensional domains, as well as integrated ensemble methods applied to approximating a real-valued function and denoising images. Performance metrics are defined in Appendix D.1, and further numerical experiments (including those in high dimensions) are provided in Appendix D.
5.1 Avoiding end-cut preference for stable cuts
To show that the ECP is remedied in MinimaxSplit, we consider a special case of the location regression model in (Cattaneo et al., 2022), similar to the discussions in Example 1. In our experiment, we generate independent replicates with $X\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1)$ and $Y$ following the standard normal or Student’s $t(\cdot)$ distributions (pure noise), sort by $X$, and scan all possible splits in the experiments in Figure 5. We also provide two baselines: Uniform random cutting picks a split point entirely at random within the current predictor range, ensuring that there is no systematic bias yet, ignoring any data-driven signal. Sampling an observed value similarly randomizes split location among actual samples. Both serve as baseline benchmarks: they eliminate ECP, but do not offer an impurity-guided improvement mechanism.
[…]